OpenStudy (anonymous):

will give metal to best answer. 78 = -2(m + 3) + m

OpenStudy (anonymous):

78 = -2(m + 3) + m 78=-2m-6+m 78+6=-m 84=-m m=-84

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw :)